# RiNALMo: General-Purpose RNA Language Models Can Generalize Well on Structure Prediction Tasks

### v1

@@ 4.2 Multi-Species Splice-Site Prediction
Identifying the donor and acceptor sites can be cast as a binary sequence-level classification task, where a method has to detect whether a query RNA sequence contains a donor/acceptor site or not. A widely used dataset of positive and negative subsets of splice-site sequences was proposed in (Scalzitti et al., 2021). The dataset was constructed by randomly selecting sequences from the exon/intron regions of the G3PO+ genomic sequences (Scalzitti et al., 2020). The dataset consists of “confirmed” error-free splice-site sequences from a diverse set of $148$ eukaryotic organisms, including humans. The test dataset consists of $4$ different species not seen in the training, namely fish (Danio rerio), fruit fly (Drosophila melanogaster), worm (Caenorhabditis elegans) and plant (Arabidopsis thaliana).
We fine-tuned RiNALMo on the training dataset to predict splice sites in the sequences. The classification token (CLS) was used as input to the prediction head. This head is configured as a two-layer multilayer perceptron (MLP) featuring a hidden layer with 128 dimensions and employing the GELU activation function. Cross-entropy loss served as our choice for the loss function. We separately fine-tuned the model, first for the donor and then for the acceptor splice-site prediction task. Finally, we compared our model’s performance with other RNA language models RNA-FM (Chen et al., 2022) and Uni-RNA (Wang et al., 2023), and several established methods such as Spliceator (Scalzitti et al., 2021) and SpliceBERT (Chen et al., 2023). We used the F1 score as the evaluation metric and presented the results in Table 3. We fine-tuned other models and used the same prediction head if they were publicly available and easy to fine-tune. Our fine-tuned model achieves state-of-the-art results in the multi-species splice-site prediction task, showing its powerful generalization properties. Notice that RiNALMo even outperforms SpliceBERT, a large language model pre-trained exclusively on pre-mRNA sequences.
Table 3: Classification F1 score for multi-species splice-site prediction. We report the average value of donor and acceptor prediction results. FT denotes whether we fine-tuned the model or represented direct citations from original papers with the same split train/test datasets.
Model FT Fish Fly Plant Worm
RiNALMo $\surd$ 97.70 96.11 96.25 95.63
SpliceBERT $\surd$ 96.70 95.83 94.96 94.82
Uni-RNA $\times$ 96.35 94.98 93.62 93.94
RNA-FM $\surd$ 94.94 93.21 89.93 90.79
Spliceator $\times$ 93.50 92.91 92.85 91.60
More details on the splice-site prediction task and the model’s hyperparameters can be found in Appendix B.2.
4.3 Mean Ribosome Loading Prediction
@@
Due to efficiency reasons, cells usually use groups of multiple ribosomes to translate the mRNA. These groups are called polyribosomes, and they enable the cell to create multiple proteins from a single mRNA molecule. To quantify protein synthesis activity, a metric called mean ribosome load (MRL), defined as the average number of ribosomes that translate the mRNA instructions into polypeptides, has been introduced. MRL can provide valuable insights into various biological functions and diseases.
@@
MRL prediction task can be viewed as a regression task where the input is the $5^{\prime}$ UTR region of the mRNA sequence. Commonly used datasets of $5^{\prime}$ UTR sequences with measured MRL values are provided by Sample et al. (2019). Two evaluation datasets were created by sampling the original dataset that contains human and random UTR sequences of varying lengths. To ensure that each sequence length is represented equally in the dataset, $100$ sequences with the deepest read coverage were selected for every length from $25$ to $100$ nucleotides. The same approach has been applied to both the human and random $5^{\prime}$ UTR sequences, yielding two evaluation datasets called Random7600 and Human7600. All remaining random $5^{\prime}$ UTRs with acceptable read coverage depth were used as the training dataset.
@@
We fine-tuned RiNALMo on the training dataset to predict the MRL value for the given $5^{\prime}$ UTR sequence. Language model outputs are fed into a prediction head which consists of six ResNet convolutional blocks. The mean squared error was used as the loss function. The model’s performance was compared to other RNA language models UNI-RNA (Wang et al., 2023) and RNA-FM (Chen et al., 2022). RNA-FM was fine-tuned with the same prediction head as RiNALMo. We also compared our model to the popular Optimus 5-prime model (Sample et al., 2019) specialized for MRL prediction. $R^{2}$ was used as the evaluation metric, and the results are reported in Table 4. Fine-tuned RiNALMo outperforms other models. Notice that RiNALMo can generalize on human UTRs despite being fine-tuned only on sequences of random origin, proving again its generalization capability.
More details on the MRL prediction task can be found in Appendix B.3.
@@
Table 4: $R^{2}$ score for mean ribosome load prediction. FT denotes whether we trained the model or represented direct citations from original papers with the same split train/test datasets.
Model FT Random7600 Human7600
RiNALMo $\surd$ 0.93 0.86
Uni-RNA $\times$ 0.91 0.85
RNA-FM $\surd$ 0.84 0.79
Optimus 5-Prime $\surd$ 0.84 0.78
@@ 5 Discussion
First, we fine-tuned and tested RiNALMo on two secondary structure prediction tasks. The first task was an intra-family secondary structure prediction, where RNAs from the same family come in both training and test datasets. The results showed that fine-tuned RiNALMo’s output representation embeds important structure information about the input sequence and, when utilized with a proper prediction head, leads to state-of-the-art performance. The second task was an inter-family secondary structure prediction, where an RNA family from the test dataset was not seen in the training dataset. It was shown previously that deep learning methods do not generalize well across RNA families (Szikszai et al., 2022). However, RiNALMo outperformed both thermodynamics-based and deep learning methods and showed that RNA language models can generalize well on unseen RNA families. This demonstrates the outstanding generalization capability of RiNALMo for secondary structure prediction tasks.
@@
Furthermore, we fine-tuned RiNALMo on two function-related RNA downstream tasks. Both of these tasks were related to the mRNA family whose examples were not contained in the pre-training dataset. The results from Section 4.2 and Section 4.3 showed that RiNALMo again generalizes well and can capture important functional information from RNA sequences from previously unseen families.
@@
In future work, we will employ RiNALMo in several other structure and function-related tasks. Of particular interest would be to test RiNALMo on tertiary structure prediction tasks and see whether its expressive output embeddings and generalization capability can improve the performance of existing or new prediction tools.
6 Conclusion
@@
In this paper, we presented RiNALMo, a new largest-to-date general-purpose RNA language model pre-trained on a dataset of $36$M unique ncRNA sequences using masked language modeling. We showed that by pre-training on a carefully curated RNA dataset and using the most modern Transformer modules and techniques, a language model can capture hidden knowledge and important structural information from unlabeled RNA sequences. The results of several structural and functional downstream tasks proved the generalization capability of RiNALMo and the expressiveness of its output representations. Of particular importance are the results of the inter-family secondary structure prediction task, where we showed that RiNALMo can generalize well on RNA families unseen during fine-tuning, unlike other deep learning methods. Due to the significance of the results, we believe RiNALMo presents a valuable asset for advancing our understanding of RNA structures and functions.

### v2

@@ 2 Results
Identifying the splice sites can be cast as a binary sequence-level classification task. A widely used dataset of positive and negative subsets of splice-site sequences was proposed in [40]. The dataset was constructed by randomly selecting sequences from the exon/intron regions of the G3PO+ genomic sequences [41]. The dataset consists of error-free splice-site sequences from a diverse set of $148$ eukaryotic organisms, including humans. The test dataset consists of four different species not seen in the training dataset.
We separately fine-tuned the model, first for donor and then for acceptor splice-site prediction. The splice-site prediction pipeline using RiNALMo embeddings is illustrated in Fig. 5b. Finally, we compared our model’s performance with other RNA LMs RNA-FM [20] and Uni-RNA [21], and several established methods such as Spliceator [40] and SpliceBERT [25]. We present the results in Fig. 5.
Figure 5: RNA splice-site prediction. a, A pre-mRNA transcript consists of non-coding, i.e., introns, and coding regions, i.e., exons. Introns are located between two exons of a gene. As part of the RNA processing pathway, introns are removed by cleavage at splice sites. These sites are found at $5^{\prime}$ and $3^{\prime}$ ends of introns, known as donor and acceptor splice sites, respectively. Most frequently, the $5^{\prime}$ end of introns begins with the dinucleotide GU, and the $3^{\prime}$ end of introns ends with AG. b, An input to RiNALMo is a $400$ nucleotide long RNA sequence from the GS_1 dataset. RiNALMo outputs $402$ embeddings including the CLS embedding. We utilize only the CLS embedding that then passes through a two-layer MLP prediction head featuring a hidden layer with $128$ dimensions and employing the GELU activation function. The output layer gives information on whether a sequence contains a donor/acceptor site or not. c, Classification F1 score for donor splice-site prediction. d, Classification F1 score for acceptor splice-site prediction. e, Classification F1 score for splice-site prediction. Here, we report the average value of donor and acceptor prediction results. In c, d and e, FT denotes whether we fine-tuned the model or represented direct citations from original papers with the same split train/test datasets. “-” denotes that the original paper did not report the result.
We fine-tuned other models and used the same prediction head if they were publicly available and easy to fine-tune. Our fine-tuned model outperforms other models, showing its powerful generalization properties. Notice that RiNALMo even outperforms SpliceBERT, an LLM pre-trained exclusively on pre-mRNA sequences. More details on the splice-site prediction task and the model’s hyperparameters can be found in Methods and Supplementary materials.
[…]
@@
Mean ribosome loading prediction using fine-tuned RiNALMo Due to efficiency reasons, cells usually use groups of multiple ribosomes to translate the mRNA. These groups are called polyribosomes, enabling the cell to create multiple proteins from a single mRNA. To quantify protein synthesis activity, an MRL metric, defined as the average number of ribosomes that translate the mRNA instructions into polypeptides, has been introduced.
@@
MRL prediction task can be viewed as a regression task where the input is the $5^{\prime}$ UTR region of the mRNA. Commonly used datasets of $5^{\prime}$ UTR sequences with measured MRL values are provided by Sample et al. [42]. Two evaluation datasets, namely Random7600 and Human7600, were created by sampling the original dataset containing human and random UTR sequences.
@@
We fine-tuned RiNALMo to predict the MRL value for the given $5^{\prime}$ UTR sequence. The prediction pipeline is illustrated in Fig. 6b. RiNALMo outputs are fed into a prediction head consisting of six ResNet blocks. The mean squared error was used as the loss function. The model’s performance was compared to other RNA LMs Uni-RNA [21] and RNA-FM [20]. We also compared our model to the popular Optimus 5-prime model [42] specialized for MRL prediction. $R^{2}$ was used as the evaluation metric, and the results are reported in Fig. 6c.
@@
Figure 6: Mean ribosome loading prediction. a, To produce multiple proteins from a single mRNA strand, cells utilize multiple ribosomes. The mean ribosome load (MRL) metric was designed to indicate the translation efficiency of a certain mRNA. b, Output embeddings produced by RiNALMo are forwarded to the one-dimensional convolutional ResNet which outputs the mean ribosome load prediction. c, $R^{2}$ score for MRL prediction on the Random7600 and Human7600 datasets. FT denotes whether we trained the model or represented direct citations from the original papers with the same split train/test datasets.
Fine-tuned RiNALMo outperforms other models. Notice that RiNALMo can generalize on human UTRs despite being fine-tuned only on sequences of random origin, again proving its generalization capability. More details on MRL prediction can be found in Methods and Supplementary materials.
@@ 3 Discussion
First, we fine-tuned and tested RiNALMo on two secondary structure prediction tasks. The first task was an intra-family secondary structure prediction, where RNAs from the same family come in both training and test datasets. The results showed that fine-tuned RiNALMo’s output representation embeds important structure information about the input sequence and, when utilized with a proper prediction head, leads to state-of-the-art performance. The second task was an inter-family secondary structure prediction, where an RNA family from the test dataset was not seen in the training dataset. It was shown previously that deep learning methods do not generalize well across RNA families [38]. However, RiNALMo outperformed both thermodynamics-based and deep learning methods and showed that RNA language models can generalize well on unseen RNA families. This demonstrates the outstanding generalization capability of RiNALMo for secondary structure prediction tasks.
@@
Furthermore, we fine-tuned RiNALMo on two function-related RNA downstream tasks. Both of these tasks were related to the mRNA family whose examples were not contained in the pre-training dataset. The results from Section 2 showed that RiNALMo again generalizes well and can capture important functional information from RNA sequences from previously unseen RNA types.
@@
In future work, we will explore the augmentation of the pre-training dataset by adding coding RNAs and evaluate how it affects the model performance on the structural and functional downstream tasks. Having bigger data might require a larger RNA LM which we also have in consideration. Multimodal pre-training including sequence and chemical mapping data is another direction for improving the representations for structure-related downstream tasks which we will explore in the future. Finally, we plan to employ RiNALMo in several other structure and function-related tasks. Of particular interest would be to test RiNALMo on tertiary structure prediction tasks and see whether its expressive output embeddings and generalization capability can improve the performance of existing or new prediction tools. Another interesting application would be employing RiNALMo’s embeddings for sequence conditioning for RNA design [43].
@@
To conclude, we presented RiNALMo, a new largest-to-date general-purpose RNA LM pre-trained on a dataset of $36$M ncRNA sequences using MLM. We showed that by pre-training on a carefully curated RNA dataset and using the most modern Transformer techniques, an LM can capture hidden knowledge and important structural information from unlabeled RNA sequences. The results on downstream tasks proved the generalization capability of RiNALMo and the expressiveness of its output representations. Of particular importance are the results of the inter-family secondary structure prediction task, where we showed that RiNALMo can generalize well on RNA families unseen during fine-tuning, unlike other DL methods. Due to the significance of the results, we believe RiNALMo presents a valuable asset for advancing our understanding of RNA structures and functions.
4 Methods
4.1 RNA Language Model
RiNALMo is an encoder-only Transformer. An input RNA sequence is tokenized, turned into a $1280$ dimension vector using a learned input embedding model and passed to the Transformer. RiNALMo consists of $33$ Transformer blocks, and each block comprises a multi-head attention and a feed-forward network (FFN). The position of the tokens is encoded using the RoPE [29], which effectively encapsulates both the relative and the absolute positional information. Each multi-head attention has $20$ attention heads. The multi-head attention is illustrated in Supplementary materials. To improve the pre-training efficiency of such a large model, we employed the IO-aware FlashAttention-2 [31], a fast and memory-efficient exact attention. In the FFN, we use two linear layers and the SwiGLU activation function [30] which combines the advantages of the Swish activation function and the gated linear unit (GLU). The FFN layers have hidden size $d_{ff}=3413$, scaling the model to $650$M parameters. The Transformer modules are interconnected using residual connections. We use the layer normalization put inside the residual blocks to stabilize the training and have well-behaved gradients at initialization [44]. We did an ablation study for RoPE and the SwiGLU activation function which showed the effectiveness of these techniques on the downstream tasks compared to the conventional absolute positional encoding and GELU activation function used in RNA-FM. The ablation study results are provided in Supplementary materials.
4.2 Tokenization
In this work, each nucleotide is a single token. During tokenization, we replace all “U”s in the sequences with “T”s. This leads to a vocabulary involving four main types of nucleotides (“A”, “C”, “T” and “G”) and several other types of combinations (“I”, “R”, “Y”, “K”, “M”, “S”, “W”, “B”, “D”, “H”, “V”, “N”, and “-”). We use additional tokens used commonly in MLM, such as [CLS], [EOS], [PAD], and [MASK]. During masking, we change nucleotides with the [MASK] token or replace it with one of the four main types of nucleotides from the vocabulary or the “any nucleotide” (“N”) token. Tokens [CLS] and [EOS] are added at the beginning and end of the sequence. The [PAD] token is appended at the end of shorter sequences to have all the sequences in a batch of the same length.
4.3 Data Preprocessing
For pre-training dataset preprocessing, we implemented a multi-step preparation pipeline. First, we collected non-coding RNA sequences from publicly available datasets RNAcentral, nt, Rfam and Ensembl. We removed sequences shorter than $16$ and longer than $8192$. Furthermore, we also removed sequence duplicates with seqkit rmdup and the resulting unique sequences were clustered with mmseqs easy-linclust with options –min-seq-id $0.7$ and -c $0.8$. Finally, we saved the processed dataset into LMDB (Lightning Memory Mapped Database) so we could easily and quickly access any data sample during the pre-training. In the end, the dataset consisted of $36$M unique ncRNA sequences clustered into $17$M clusters. In comparison, Chen et al. [20] collected ncRNAs from the RNAcentral database only. As a preprocessing step, the authors removed only duplicate RNA sequences. We clustered the RNA sequences to ensure sequence diversity in each training batch since we later sample from each cluster once per epoch.
4.4 Pre-training
We pre-trained RiNALMo using the MLM task where we corrupted unlabeled RNA sequences and then tasked the model to reconstruct them. To corrupt the input sequence, we randomly selected $15\%$ of the tokens in the training sequence. Of these, $80\%$ are masked, i.e., replaced with the unique vocabulary token [MASK], $10\%$ are replaced with a randomly selected token from the vocabulary, and the remaining $10\%$ are left intact.
Let us denote the corrupted sequence by ${\tilde{\bm{X}}=\{\tilde{x}_{i}\}}$. We train RiNALMo parameterized by $\theta$ to reconstruct ${\bm{X}}$ by predicting the masked tokens conditioned on $\tilde{\bm{X}}$. For a given input token $\tilde{x}_{i}$, the loss is the probability of the correct nucleotide, given $\tilde{\bm{X}}$:
$\mathcal{L}_{MLM}=-\log p_{\theta}(\tilde{x}_{i}=x_{i}|\tilde{\bm{X}}).$
where $\mathcal{M}$ is the index set of masked tokens.
Due to the computational limitations, during training, we limited the length of the input into the language model to $1024$ tokens. Every sequence begins with a [CLS] token, i.e., a classification token, and ends with an [EOS] token, i.e., an end-of-sequence token. If an input sequence is longer than $1022$ nucleotides, a random part of the sequence will be cropped out and fed into the model during each epoch, similar to what was done in [13]. To implicitly give information to the model that the sequence was cropped, it does not begin with the [CLS] token nor ends with the [EOS] token unless we crop the beginning or end of the sequence.
To ensure sequence diversity in each training batch during pre-training, we randomly sampled each sequence in the batch from a different sequence cluster. Effectively, it means that in every epoch RiNALMo saw $17$M samples, i.e., one sequence from each of the clusters, and in each epoch, sequences were sampled randomly from the clusters using a new seed.
We pre-trained RiNALMo with seven A100 GPUs of $80$ GB memory for two weeks. The batch size was set to $192$ per GPU. We adopted the cosine annealing learning rate schedule with a linear warm-up. During the warm-up period, learning rate increases from $10^{-7}$ to $5{\times}10^{-5}$ for $2000$ steps. For the cosine annealing schedule, the minimum learning rate was set to $10^{-5}$. The gradient norm was clipped to 1.0.
We pre-trained several configurations of the LM: RiNALMo-$650$M with $650$M parameters, RiNALMo-$150$M with $148$M parameters, and RiNALMo-$33$M with $33.5$M parameters. We did several experiments to evaluate how the model’s size influences its performance on the pre-training and downstream tasks. First, we measured perplexity. Perplexity is a metric used to evaluate how well an LM predicts a sample of text or, in our case, an RNA sequence. The perplexity is defined as the exponential of the negative log-likelihood of the sequence. To efficiently calculate the perplexity we used the following term
$\textsc{Perplexity}(x)=\exp\{-\log p_{\theta}(\tilde{x}_{i\in\mathcal{M}}=x_{i}|\tilde{\bm{X}})\}.$
Lower perplexity means the LM is better at reconstructing the masked tokens in the MLM pre-training task.
In order to evaluate how the number of parameters affects the model’s ability to reconstruct the masked tokens, we calculated the perplexity of RiNALMo for several different model sizes. Validation perplexity was measured on a 1% random-split holdout of the pre-training dataset. The perplexity values are given in Extended Data Fig. 2. It can be seen that as we increase the model size, the perplexity decreases, meaning the larger the LM gets, it is more capable of reconstructing the missing tokens.
The hyperparameters of RiNALMo models of different sizes and the comparison of their performance on the downstream tasks are given in Supplementary materials. Based on these results we can conclude that besides novel techniques such as RoPE and SwiGLU, model size is important in achieving a boost in performance on pre-training and consequently downstream tasks.
4.5 Secondary Structure Prediction
To adapt RiNALMo for the secondary structure prediction, we fine-tuned the model with a simple binary classification task with binary cross-entropy loss, where we tasked the model to classify each nucleotide pair as either paired or unpaired. A simple bottleneck residual neural network (ResNet) [45] is attached to the language model and is fed with nucleotide pair representations. These representations are obtained by applying outer concatenation to RiNALMo’s output. For example, to obtain the vector representation of the nucleotide pair $(i,j)$, we simply concatenate the representation of the nucleotide $j$ to the representation of the nucleotide $i$.
In the end, our secondary structure prediction pipeline outputs a matrix where each element represents a pairing probability logit for a certain nucleotide pair. Because of the symmetry of secondary structures (if nucleotide $i$ is paired with nucleotide $j$, then $j$ is paired with $i$ as well), we calculate training loss only on matrix elements “above” the main diagonal.
During fine-tuning, we utilized gradual parameter unfreezing. After every three epochs, we unfroze additional three RiNALMo’s layers. In the first three epochs, we trained only the prediction head. The model was fine-tuned for $15$ epochs and the learning rate was initially set to $10^{-4}$ with a linear decay schedule. We used the same prediction head architecture for RNA-FM fine-tuning. Due to architectural differences between RiNALMo and RNA-FM, we decided to modify the fine-tuning schedule. The prediction head was first pre-trained while RNA-FM parameters were kept frozen. After three epochs we unfroze RNA-FM and fine-tuned it alongside the prediction head.
When assessing the performance of secondary structure prediction, it is important to consider RNA structural dynamics, and that’s why it is helpful to view predictions that are close enough as correct. We employed the metric calculation approach proposed in [46]. To be more precise, for a nucleotide pairing $(i,j)$ where $i$ and $j$ represent nucleotide indices in the RNA sequence, $(i\pm 1,j)$ and $(i,j\pm 1)$ pairings are also considered correct predictions. To obtain the F1 scores we reported in this study, we calculated the F1 score separately on each structure and then averaged those values.
In Supplementary materials, we provide the performance of smaller LMs, the ablation study and the impact of fine-tuning. We also report sequence-length weighted F1 scores on the inter-family secondary structure prediction task.
4.6 Multi-Species Splice-Site Prediction
Splice-site prediction task is essentially a binary sequence-level classification task, where a method has to detect whether a query RNA sequence contains a donor/acceptor site or not. We fine-tuned RiNALMo on the training dataset to predict splice sites in the sequences. The classification token (CLS) was used as input to the prediction head. This head is configured as a two-layer multilayer perceptron (MLP) featuring a hidden layer with 128 dimensions and employing the GELU activation function. Cross-entropy loss served as our choice for the loss function. We separately fine-tuned the model, first for the donor and then for the acceptor splice-site prediction task.
We used the dataset denoted as GS_1 in the paper it was proposed in Scalzitti et al. [40]. GS_1 has the same ratio of positive to negative samples, whose negative samples consist of exon, intron, or false positive sequences. The length of the input sequences was set to $400$ nucleotides for both the training and test datasets.
The prediction head was uniformly initialized from $\mathcal{U}(-\sqrt{1/d},\sqrt{1/d})$, where $d$ is the input features dimension. The learning rate was set to $10^{-5}$ and the language model with the prediction head was fine-tuned for $2$ epochs when it achieved the best results on the validation dataset. The batch size was set to $32$.
We fine-tuned the SpliceBERT model and the RNA-FM model in combination with a two-layer MLP prediction head. We fine-tuned the models on the same training/validation dataset used for fine-tuning RiNALMo with the same learning rate.
More splice-site prediction results can be found in Supplementary materials.
4.7 Mean Ribosome Loading Prediction
We fine-tuned RiNALMo on the appropriate training dataset to predict the MRL value for the given $5^{\prime}$ UTR sequence. Language model outputs are fed into a prediction head which consists of six ResNet convolutional blocks. The mean squared error was used as the loss function. RNA-FM was fine-tuned with the same prediction head as RiNALMo.
Each block of the prediction head consists of two 1D convolution layers followed by instance normalization and ELU activation function. Parameters of the prediction head were initialized with the default Pytorch [47] parameter initialization method.
Training and evaluation datasets were produced with the procedure described in [42]. Two evaluation datasets were created by sampling the original dataset that contains human and random UTR sequences of varying lengths. To ensure that each sequence length is represented equally in the dataset, $100$ sequences with the deepest read coverage were selected for every length from $25$ to $100$ nucleotides. The same approach has been applied to both the human and random $5^{\prime}$ UTR sequences, yielding two evaluation datasets called Random7600 and Human7600. All remaining random $5^{\prime}$ UTRs with acceptable read coverage depth were used as the training dataset.
More MRL prediction results can be found in Supplementary materials.
5 Data availability
We used several databases with unannotated RNA sequences, namely RNAcentral [23], nt [48], Rfam [49] and Ensembl [50]. The RNAcentral dataset is available at https://ftp.ebi.ac.uk/pub/databases/RNAcentral/current_release/sequences/. The nt dataset is available at https://ftp.ncbi.nlm.nih.gov/blast/db/FASTA/. The Rfam dataset is available at https://ftp.ebi.ac.uk/pub/databases/Rfam/CURRENT/fasta-files/Rfam.fa. Finally, the Ensembl dataset is available at https://ftp.ensembl.org/pub/release-109/fasta/.
The intra-family RNA secondary structure dataset with train/test splits is available at https://dl.dropboxusercontent.com/s/w3kc4iro8ztbf3m/bpRNA_dataset.zip. The inter-family secondary structure dataset consisting nine families and train/test splits is available at https://github.com/marcellszi/dl-rna/releases/download/Data/ct-splits.tar.gz.
For the multi-species splice-site downstream task, we used the dataset denoted as GS_1 in [40]. The dataset is available at https://git.unistra.fr/nscalzitti/spliceator and https://zenodo.org/records/7995778.
The mean ribosome loading dataset [42] used in the paper can be found at https://www.ncbi.nlm.nih.gov/geo/download/?acc=GSE114002&format=file.
6 Code availability
The code repository is available on https://github.com/lbcb-sci/RiNALMo and the pre-trained and fine-tuned weights are available on https://zenodo.org/records/10725749. The weights can be automatically downloaded using the script provided in the repository. We provide scripts for fine-tuning the pre-trained model on the downstream tasks from the Results section. The data used in the downstream tasks can be automatically downloaded and preprocessed using the scripts in the code repository.